​ In the Keynesian cross, assume that the consumption function is given by

 In the Keynesian cross, assume that the consumption function is given by 

C=200+0.75 (Y-T) 

Planned investment is 100; government purchases and taxes are both 100. 

a. Graph planned expenditure as a function of income. 

b. What is the equilibrium level of income?
Answer #1

Y= (t ta AD=Y= 200 +0.75(4-100) + 100 + 100 AD= 200 +0.75Y - 75 +200 AD= 400-75to.ASY AD= 325 +0.754 equilibican Y=) Y= 325to3zsto.try Planned expenditure Output